# Coexistence of Chiral Majorana Edge States and Bogoliubov Fermi Surfaces in Two-Dimensional Nonsymmorphic Dirac Semimetal/Superconductor Heterostructures

### v1

@@ Abstract
Dirac semimetals are renowned for the host of singular symmetry-protected band degeneracies which can give rise to other exotic phases. In this work, we consider a two-dimensional Dirac semimetal stabilized by $\mathcal{PT}$ symmetry and nonsymmorphic symmetries. We find that an out-of-plane Zeeman field can lift the Dirac points and transform the system into a Chern insulator with chiral edge states. By placing the nonsymmorphic Dirac semimetal in proximity to an $s$-wave superconductor, we uncover that chiral topological superconductors with large Chern numbers can be achieved. In addition, we find that topologically-protected Bogoliubov Fermi surface can also emerge in this system, due to the coexistence of inversion symmetry and particle-hole symmetry. Notably, we find that the chiral Majorana edge state persists even when the Chern number becomes ill-defined due to the appearance of Bogoliubov Fermi surfaces. The impact of these Bogoliubov Fermi surfaces on the thermal Hall effects is also investigated. Our study not only identifies a class of materials capable of realizing topological Bogoliubov Fermi surfaces through conventional $s$-wave superconductivity, but also uncovers an exotic phase where chiral Majorana edge states and Bogoliubov Fermi surfaces coexist.
@@
In the past two decades, chiral topological superconductors (CTSCs) have garnered significant attention due to the host of many exotic properties[1, 2, 3, 4, 5, 6, 7, 8]. CTSCs refer to two-dimensional (2D) fully-gapped superconductors characterized by a nonzero Chern number. Their defining characteristics include the presence of robust chiral Majorana edge states and quantized thermal Hall effects in the low-temperature limit[9, 10]. Additionally, when the Chern number is odd, vortices in the CTSCs will harbor robust Majorana zero modes which exhibit non-Abelian statistics[11] and hold potential for applications in topological quantum computation[12, 13, 14]. While CTSCs can naturally arise from pairings with nonzero angular momentum, such as $p+ip$, $d+id$ and $f+if$ pairings[15, 16, 17], the scarcity of such pairings in superconducting materials has prompted the exploration of alternative pathways. A significant breakthrough in theory is the recognition that the combination of antisymmetric spin-orbit coupling (SOC), Zeeman field with direction perpendicular to the spin plane of the SOC, and conventional $s$-wave pairing-all prevalent in real materials-can lead to the realization of CTSCs[18, 19, 20, 21, 22, 23, 24, 25]. This theoretical advancement has sparked experimental efforts to pursue CTSCs, and notable progress has been witnessed in this endeavor[26, 27, 28, 29].
@@
Recently, Bogoliubov Fermi surfaces (BFSs) in superconductors have generated considerable interest due to their profound impact on various transport properties of these materials[30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. BFSs represent the Fermi surfaces of Bogoliubov quasiparticles, signifying a finite density of states for zero-energy excitations. Typically, the presence of zero-energy excitations is detrimental to superconductors, as it suggests that breaking Cooper pairs incurs no energy cost. However, recent theoretical studies have demonstrated that BFSs can be stabilized under specific conditions[46, 47, 48, 49, 50]. Intriguingly, while a global topological invariant defined over the Brillouin zone is no longer viable for superconductors with BFSs, the BFS itself can possess topological protection[51, 52]. In a seminal paper, Agterberg et al. revealed that multiband systems with intrinsic time-reversal-symmetry-breaking interband pairings can exhibit BFSs protected by a $Z_{2}$ invariant if the inversion symmetry is preserved[53]. In the study of pairing in 3D spin-orbit coupled spin-$3/2$ semimetals, it is also discovered that some unconventional pairings can lead to BFSs[54, 55, 56, 57, 58]. Besides BFSs induced by pairings breaking time-reversal symmetry intrinsically, Yuan and Fu also revealed that the combination of antisymmetric SOC, Zeeman field with direction lying in the spin plane of the SOC, and conventional $s$-wave pairing can also give rise to BFSs[59]. This scenario was later demonstrated in superconductor-topological insulator [60] and superconductor-semiconductor hybrid systems[61].
@@
Interestingly, the combination of antisymmetric SOC, Zeeman field, and conventional $s$-wave pairing can give rise to both CTSCs and BFSs in two dimensions. However, these two phenomena have distinct requirements: CTSCs necessitate a perpendicular Zeeman field[21, 22], whereas BFSs require a parallel Zeeman field aligned with the spin plane of the SOC[59]. Furthermore, it is important to note that since the Chern number is undefined in a gapless superconductor, CTSCs and BFSs inherently represent two incompatible phenomena. Therefore, it appears plausible to conclude that BFSs and chiral Majorana edge states would typically exist exclusively under their respective conditions. In this work, we investigate a 2D nonsymmorphic spin-orbit coupled Dirac semimetal (DSM) under the influence of a conventional $s$-wave superconductor and the effect of a perpendicular Zeeman field, as illustrated in Fig.1(a). Our findings demonstrate that, contrary to expectations, BFSs and chiral Majorana edge states can coexist within this system.
@@
Figure 1: (a) Schematic diagram of the system. (b) The lattice structure of the DSM. (c) The Brillouin zone and the four time-reversal invariant momentums. (d) Band structure at $t=0$, $\lambda=0.5$, $\eta=0.5$ and $M_{z}=0$. (e) Band structure at $t=0$, $\lambda=0.5$, $\eta=0.5$ and $M_{z}=0.3$.
@@
2D nonsymmorphic DSM and Zeeman-field-induced Chern insulator.—In two dimensions, stabilizing fourfold degenerate Dirac points in the presence of SOC requires not only $\mathcal{PT}$ symmetry, but also nonsymmorphic symmetries[62]. In this work, we consider a DSM realized on the bipartite checkerboard lattice as illustrated in Fig.1(b). The tight-binding Hamiltonian describing the DSM is given by[62]
@@
under the basis $\psi^{{\dagger}}_{\boldsymbol{k}}=(c_{A,\uparrow,\boldsymbol{k}}^{{\dagger}},c_{B,\uparrow,\boldsymbol{k}}^{{\dagger}},c_{A,\downarrow,\boldsymbol{k}}^{{\dagger}},c_{B,\downarrow,\boldsymbol{k}}^{{\dagger}})$. Here $\epsilon_{0}(\boldsymbol{k})=-2t(\cos k_{x}+\cos k_{y})$ denotes the kinetic energy arising from the nearest-neighbor intra-sbulattice hoppings, $\epsilon_{1}(\boldsymbol{k})=-4\eta\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ denotes the kinetic energy originating from the nearest-neighbor inter-sublattice hoppings, and $\Lambda_{\rm so}(\boldsymbol{k})=2\lambda_{\rm so}(\sin k_{y}s_{x}-\sin k_{x}s_{y})$ is the SOC. The Pauli matrices $\sigma_{i}$ and $s_{i}$ act on the sublattice ($A,B$) and spin ($\uparrow$,$\downarrow$ , ↓) degrees of freedom, respectively. For notational simplicity, the lattice constants are set to unity throughout this work, and identity matrices in orbital and spin space are made implicit.
The Hamiltonian in Eq.(1) has inversion symmetry ($\mathcal{P}=\sigma_{x}$) despite the presence of SOC. Such inversion-symmetric SOC can emerge when the system locally breaks the inversion symmetry but preserve it globally[63]. The Hamiltonian also has time-reversal symmetry ($\mathcal{T}=is_{y}\mathcal{K}$ with $\mathcal{K}$ the complex conjugation operator), glide symmetry $\{\mathcal{M}_{z}|(\frac{1}{2},\frac{1}{2})\}$, and two screw symmetries $\{\mathcal{C}_{2x}|(\frac{1}{2},0)\}$ and $\{\mathcal{C}_{2y}|(0,\frac{1}{2})\}$[62]. Because of the $\mathcal{P}\mathcal{T}$ symmetry and the three nonsymmorphic symmetries, this Hamiltonian harbors three fourfold degenerate Dirac points at the boundary of the Brillouin zone, with their positions located at the three time-reversal invariant momentums (TRIMs), $\mathbf{X}=(\pi,0)$, $\mathbf{Y}=(0,\pi)$ and $\mathbf{M}=(\pi,\pi)$, as shown in Figs.1(c) and 1(d).
It is known that breaking time-reversal symmetry can lift Dirac points, leading to Chern insulators with gapless chiral edge states[64, 65, 66]. The most straightforward way to break the time-reversal symmetry is by introducing a spin-splitting Zeeman field. However, for DSMs that lack SOC and thus possess full spin-rotation symmetry, such as monolayer graphene, a Zeeman field is unable to gap the Dirac points. Instead, it will shift the bands with opposite spin polarizations in opposite directions in energy, resulting in the splitting of fourfold spin-degenerate Dirac points into twofold spin-polarized Dirac points. For the spin-orbit coupled DSM described by the Hamiltonian in Eq.1, we find that the situation is remarkably different. To illustrate this, consider an out-of-plane Zeeman field, where the field direction is perpendicular to the spin plane of the SOC. Accordingly, the Hamiltonian becomes
@@
where $M_{z}=g\mu_{B}B_{z}/2$ denotes the strength of Zeeman field induced, e.g., by a $z$-directional magnetic field, i.e., $\mathbf{B}=B_{z}\hat{\mathbf{z}}$. As the first term does not affect the band topology, in this part we let $\epsilon_{0}(\boldsymbol{k})=0$ to simplify the discussion. The corresponding energy spectra are given by
@@
where $\Lambda(\boldsymbol{k})=2\lambda_{\rm so}\sqrt{\sin^{2}k_{x}+\sin^{2}k_{y}}$, and $\alpha=\pm 1$ and $\beta=\pm 1$. It is easy to see that, once $M_{z}\neq 0$, the three Dirac points at $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{M}$ are all gapped, as depicted in Fig.1(e). However, as $M_{z}$ reaches ±$\pm 4\eta$, the energy gap closes and reopens at $\boldsymbol{\Gamma}=(0,0)$, indicating the occurrence of topological phase transitions. Since the energy gap remains open whenever the absolute value of $M_{z}$ exceeds $4\eta$ (assuming $\eta$ is positive), it is immediately evident that the system is topologically trivial in this regime. Conversely, in the range of $0<|M_{z}|<4\eta$, the gapped phases are expected to be topological. To confirm this, we numerically calculate the Chern number and determine the phase diagram shown in Fig.2(a). Within the topological region, we find that the system with a cylindrical geometry possesses one gapless chiral edge state, which aligns with the calculated Chern number, as illustrated in Fig.2(b).
Figure 2: (a) Phase diagram valid as long as $\lambda_{\rm so}\neq 0$. (b) A Chern insulator with one branch of chiral edge state. The parameters are $t=0$, $\lambda_{\rm so}=0.5$, $\eta=0.5$ and $M_{z}=0.3$.
Nonsymmorphic DSM in proximity to an $s$-wave superconductor.—The SOC in the DSM holds the promise of topological superconductivity if superconductivity is successfully induced into the DSM. In this paper, we explore the scenario where the DSM is grown on the surface of a conventional fully-gapped $s$-wave superconductor [see Fig.1(a)]. Through the proximity effect, the DSM inherits $s$-wave pairing from the underlying superconductor. The superconducting DSM is effectively described by the Bogoliubov-de Gennes (BdG) Hamiltonian which is of the form
@@
Without the Zeeman field, the BdG Hamiltonian falls into the DIII class according to the tenfold-way classification[68, 69]. Due to the momentum-independent nature of $s$-wave pairing, the superconducting phase is always topologically trivial in this case. However, upon breaking time-reversal symmetry through the application of a Zeeman field, the Hamiltonian transitions to the D class, thereby enabling the existence of CTSCs. Regardless of the presence or absence of the Zeeman field, the BdG Hamiltonian possesses both inversion symmetry and particle-hole symmetry ($\Xi=\tau_{x}\mathcal{K}$), resulting in a symmetric energy spectrum centered at $E=0$. These two symmetries are crucial for the topological protection of the BFSs discussed later.
To demonstrate that the combination of Dirac band structure, a Zeeman field, and $s$-wave pairing can lead to the emergence of CTSCs, we initially consider the case where $\eta=0$. In this limit, the two sublattices decoupled, i.e., $\mathcal{H}(\boldsymbol{k})=\mathcal{H}_{A}(\boldsymbol{k})\oplus\mathcal{H}_{B}(\boldsymbol{k})$, where
@@
±$\pm 2\lambda_{\rm so}(\sin k_{x}\tau_{z}s_{y}-\sin k_{y}s_{x})+\Delta_{s}\tau_{y}s_{y}.$
The SOC simplifies to the Rashba form, and this particular form of the Hamiltonians $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ is well-known for supporting CTSCs[21, 22]. Despite the fact that the SOC has opposite signs for $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, their Chern numbers are identical. To facilitate a clear discussion, we assume that $t$, $M_{z}$, and $\Delta_{s}$ are all positive, with $\Delta_{s}$ being less than $M_{z}$. By varying only the chemical potential $\mu$, we determine the dependence of the Chern number on $\mu$[70], which is given by
@@
where $\mu_{c}=\sqrt{M_{z}^{2}-\Delta_{s}^{2}}$. Since $C_{A}$ equals $C_{B}$, the Chern number of the resulting CTSCs, $C=C_{A}+C_{B}$, will always be even. Intriguingly, the maximum value of $C$, and consequently the maximum number of chiral Majorana edge states, can reach as high as $4$.
@@
Intriguingly, we find that when $\eta$ attains a critical threshold, the bulk energy gap closes. However, further increase in $\eta$ fails to reopen the gap. Instead, the middle two bands remain in contact, forming line degeneracies. Notably, despite the absence of a bulk gap resulting in an undefined Chern number, the chiral Majorana edge states persist if their spectral crossing occurs at $k_{y}=\pi$ [$k_{x}=\pi$ if open (periodic) boundary conditions are considered in the $y$ ($x$) direction due to $C_{4z}$ rotation symmetry] of the edge Brillouin zone as illustrated in Figs.3(c) and 3(d). The resilience of the chiral Majorana edge states can be attributed to the vanishing of the $\epsilon_{1}(\boldsymbol{k})$ term when $k_{x}$ or $k_{y}$ equals $\pi$. In simpler terms, when the spectral crossing of these states occurs at $k_{x}=\pi$ or $k_{y}=\pi$, an increase in $\eta$ has no influence on the edge gap or the crossing at that momentum, as $\epsilon_{1}(\boldsymbol{k})$ identically vanishes at those points. In contrast, if the spectrum crossing occurs at $k_{x}=0$ or $k_{y}=0$, an increase in $\eta$ will cause both the edge gap and the spectrum crossing at that momentum to disappear, as illustrated in Fig.3(d).
Figure 3: Energy spectrum for a system of open (periodic) boundary conditions in the $x$ ($y$) direction. (a) TSC with $C=2$. (b) TSC with $C=-4$. Figures (c) and (d) show the coexistence of BFS and chiral Majorana edge states. All figures have common parameters $t=1$, $\lambda=0.5$, $M_{z}=0.5$ and $\Delta_{s}=0.2$. The parameters $(\mu,\eta)$ are set to $(4,0.1)$ in (a), $(0,0.025)$ in (b), $(4,0.8)$ in (c) and $(0,0.2)$ in (d), respectively.
In two dimensions, line degeneracy in a superconductor corresponds to topologically protected BFSs. As aforementioned, the topological protection originates from the coexistence of inversion symmetry and particle-hole symmetry in this system. As firstly noted in Ref.[53], when these two symmetries coexist, the Hamiltonian can be transformed into an antisymmetric form, given by ~$\tilde{\mathcal{H}}(\boldsymbol{k})=U\mathcal{H}(\boldsymbol{k})U^{{\dagger}}$, where $U$ denotes a unitary matrix. The transformed Hamiltonian obeys the relation ~$\tilde{\mathcal{H}}(\boldsymbol{k})=-\tilde{\mathcal{H}}^{\mathbb{T}}(\boldsymbol{k})$, with $\mathbb{T}$ denoting the transpose operation. The antisymmetric property allows for a definition of Pfaffian, $P(\boldsymbol{k})=\text{Pf}[\tilde{H}(\boldsymbol{k})]$. Since $\det\mathcal{H}(\boldsymbol{k})=\det\tilde{\mathcal{H}}(\boldsymbol{k})=P^{2}(\boldsymbol{k})$, the zeros of $P(\boldsymbol{k})$ determines the band degeneracies of $\mathcal{H}(\boldsymbol{k})$ at $E=0$. Topologically protected BFSs correspond to the line degeneracy across which the sign of $P(\boldsymbol{k})$, denoted as $\text{sgn}[P(\boldsymbol{k})]$, changes from $1$ to $-1$, or vice versa.
For the current Hamiltonian, $P(\boldsymbol{k})$ is given by (see Supplemental Material[71])

### v2

@@ Abstract
Dirac semimetals are renowned for the host of singular symmetry-protected band degeneracies which can give rise to other exotic phases. In this work, we consider a two-dimensional Dirac semimetal stabilized by $\mathcal{PT}$ symmetry and nonsymmorphic symmetries. We find that an out-of-plane Zeeman field can lift the Dirac points and transform the system into a Chern insulator with chiral edge states. By placing the nonsymmorphic Dirac semimetal in proximity to an $s$-wave superconductor, we uncover that chiral topological superconductors with large Chern numbers can be achieved. In addition, we find that topologically-protected Bogoliubov Fermi surfaces can also emerge in this system, due to the coexistence of inversion symmetry and particle-hole symmetry. Notably, we find that the chiral Majorana edge state persists even when the Chern number becomes ill-defined due to the appearance of Bogoliubov Fermi surfaces. The impact of these Bogoliubov Fermi surfaces on the thermal Hall effects is also investigated. Our study not only identifies a class of materials capable of realizing topological Bogoliubov Fermi surfaces through conventional $s$-wave superconductivity, but also uncovers an exotic phase where chiral Majorana edge states and Bogoliubov Fermi surfaces coexist.
@@
In the past two decades, chiral topological superconductors (CTSCs) have garnered significant attention due to the host of many exotic properties [1, 2, 3, 4, 5, 6, 7, 8, 9]. CTSCs refer to two-dimensional (2D) fully-gapped superconductors characterized by a nonzero Chern number. Their defining characteristics include the presence of robust chiral Majorana edge states and quantized thermal Hall effects in the low-temperature limit [10, 11]. Additionally, when the Chern number is odd, vortices in the CTSCs will harbor robust Majorana zero modes which exhibit non-Abelian statistics [12] and hold potential for applications in topological quantum computation [13, 14, 15, 16]. While CTSCs can naturally arise from pairings with nonzero angular momentum, such as $p+ip$, $d+id$ and $f+if$ pairings [17, 18, 19], the scarcity of such pairings in superconducting materials has prompted the exploration of alternative pathways. A significant breakthrough in theory is the recognition that the combination of antisymmetric spin-orbit coupling (SOC), Zeeman field with direction perpendicular to the spin plane of the SOC, and conventional $s$-wave pairing—all prevalent in real materials—can lead to the realization of CTSCs [20, 21, 22, 23, 24, 25, 26, 27]. This theoretical advancement has sparked experimental efforts to pursue CTSCs, and notable progress has been witnessed in this endeavor [28, 29, 30, 31].
@@
Recently, Bogoliubov Fermi surfaces (BFSs) in superconductors have generated considerable interest due to their profound impact on various transport properties of these materials [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 44]. BFSs represent the Fermi surfaces of Bogoliubov quasiparticles, signifying a finite density of states for zero-energy excitations [48]. Typically, the presence of zero-energy excitations is detrimental to superconductors, as it suggests that breaking Cooper pairs incurs no energy cost. However, recent theoretical studies have demonstrated that BFSs can be stabilized under specific conditions [49, 50, 51, 52, 53]. Intriguingly, while a global topological invariant defined over the Brillouin zone is no longer viable for superconductors with BFSs, the BFS itself can possess topological protection [54, 55]. In a seminal paper, Agterberg et al. revealed that multiband systems with intrinsic time-reversal-symmetry-breaking interband pairings can exhibit BFSs protected by a $Z_{2}$ invariant if the inversion symmetry is preserved [56]. In the study of pairing in 3D spin-orbit coupled spin-$3/2$ semimetals, it is also discovered that some unconventional pairings can lead to BFSs [57, 58, 59, 60, 61]. Besides BFSs induced by pairings breaking time-reversal symmetry intrinsically, Yuan and Fu also revealed that the combination of antisymmetric SOC, Zeeman field with direction lying in the spin plane of the SOC, and conventional $s$-wave pairing can also give rise to BFSs [62]. This scenario was later demonstrated in superconductor-topological insulator [63] and superconductor-semiconductor hybrid systems [64].
@@
Interestingly, the combination of antisymmetric SOC, Zeeman field, and conventional $s$-wave pairing can give rise to both CTSCs and BFSs in two dimensions. However, these two phenomena have distinct requirements: CTSCs necessitate a perpendicular Zeeman field [23, 24], whereas BFSs require a parallel Zeeman field aligned with the spin plane of the SOC [62]. Furthermore, it is important to note that since the Chern number is undefined in a gapless superconductor, CTSCs and BFSs inherently represent two incompatible phenomena. Therefore, it appears plausible to conclude that BFSs and chiral Majorana edge states would typically exist exclusively under their respective conditions. In this work, we investigate a 2D nonsymmorphic spin-orbit-coupled Dirac semimetal (DSM) under the influence of a conventional $s$-wave superconductor and the effect of a perpendicular Zeeman field, as illustrated in Fig.1(a). Our findings demonstrate that, contrary to expectations, BFSs and chiral Majorana edge states can coexist within this system.
@@
Figure 1: (a) Schematic of the heterostructure. (b) Lattice structure of the DSM. (c) Brillouin zone with the four time-reversal invariant momentums highlighted. (d) Band structure along high symmetry lines within the Brillouin zone. The parameter values are: $t=0$, $\lambda=0.5$, $\eta=0.5$ and $M_{z}=0$. (e) Band structure with the Zeeman field set to $M_{z}=0.3$.
@@
2D nonsymmorphic DSM and Zeeman-field-induced Chern insulator.—In two dimensions, stabilizing fourfold degenerate Dirac points in the presence of SOC requires not only $\mathcal{PT}$ symmetry, but also nonsymmorphic symmetries [65]. In this work, we consider a DSM realized on the bipartite lattice as illustrated in Fig.1(b). The tight-binding Hamiltonian describing the DSM is given by [65]
@@
under the basis $\psi^{{\dagger}}_{\boldsymbol{k}}=(c_{A,\uparrow,\boldsymbol{k}}^{{\dagger}},c_{A,\downarrow,\boldsymbol{k}}^{{\dagger}},c_{B,\uparrow,\boldsymbol{k}}^{{\dagger}},c_{B,\downarrow,\boldsymbol{k}}^{{\dagger}})$. Here $\epsilon_{0}(\boldsymbol{k})=-2t(\cos k_{x}+\cos k_{y})$ denotes the kinetic energy arising from the nearest-neighbor intra-sublattice hopping, $\epsilon_{1}(\boldsymbol{k})=-4\eta\cos\frac{k_{x}}{2}\cos\frac{k_{y}}{2}$ denotes the kinetic energy originating from the nearest-neighbor inter-sublattice hopping, and $\Lambda_{\rm so}(\boldsymbol{k})=2\lambda_{\rm so}(\sin k_{y}s_{x}-\sin k_{x}s_{y})$ is the SOC. The Pauli matrices $\sigma_{i}$ and $s_{i}$ act on the sublattice ($A,B$) and spin ($\uparrow,\downarrow$) degrees of freedom, respectively. For notational simplicity, the lattice constants are set to unity throughout this work, and identity matrices in orbital and spin space are made implicit.
The Hamiltonian in Eq.(1) has inversion symmetry ($\mathcal{P}=\sigma_{x}$) despite the presence of SOC. Such inversion-symmetric SOC can emerge when the system locally breaks the inversion symmetry but preserves it globally [66]. The Hamiltonian also has time-reversal symmetry ($\mathcal{T}=is_{y}\mathcal{K}$ with $\mathcal{K}$ the complex conjugate operator), a glide mirror symmetry $\{\mathcal{M}_{z}|(\frac{1}{2},\frac{1}{2})\}$, and two screw symmetries $\{\mathcal{C}_{2x}|(\frac{1}{2},0)\}$ and $\{\mathcal{C}_{2y}|(0,\frac{1}{2})\}$[65]. Because of the $\mathcal{P}\mathcal{T}$ symmetry and the three nonsymmorphic symmetries, this Hamiltonian harbors three fourfold degenerate Dirac points at the boundary of the Brillouin zone, with their positions located at the three time-reversal invariant momentums (TRIMs), $\mathbf{X}=(\pi,0)$, $\mathbf{Y}=(0,\pi)$ and $\mathbf{M}=(\pi,\pi)$, as shown in Figs.1(c) and 1(d).
It is known that breaking time-reversal symmetry can lift Dirac points, resulting in Chern insulators characterized by gapless chiral edge states [67, 68, 69]. The most straightforward way to break time-reversal symmetry is through the application of a spin-splitting Zeeman field. However, in DSMs lacking SOC and thus exhibiting full spin-rotation symmetry (e.g., monolayer graphene), a Zeeman field cannot open a gap at the Dirac points. Instead, it shifts the energy bands with opposite spin polarizations in opposite directions, causing the fourfold spin-degenerate Dirac points to split into twofold spin-polarized Weyl points. In contrast, for the spin-orbit-coupled DSM described by the Hamiltonian in Eq.1, we observe a markedly different behavior. To demonstrate this, consider an out-of-plane Zeeman field oriented perpendicular to the spin plane of the SOC. Accordingly, the Hamiltonian becomes
@@
where $M_{z}=g\mu_{B}B_{z}/2$ denotes the strength of Zeeman field induced, e.g., by a $z$-directional magnetic field, i.e., $\mathbf{B}=B_{z}\hat{\mathbf{z}}$. As the first term does not affect the band topology, in this part we let $\epsilon_{0}(\boldsymbol{k})=0$ by setting $t=0$ to simplify the discussion. The corresponding energy spectra are given by
@@
where $\Lambda(\boldsymbol{k})=2\lambda_{\rm so}\sqrt{\sin^{2}k_{x}+\sin^{2}k_{y}}$, $\alpha=\pm 1$ and $\beta=\pm 1$. It is easy to see that, once $M_{z}\neq 0$, the three Dirac points at $\mathbf{X}$, $\mathbf{Y}$ and $\mathbf{M}$ are all gapped, as depicted in Fig.1(e). However, as $M_{z}$ reaches $\pm 4\eta$, the energy gap closes and reopens at $\boldsymbol{\Gamma}=(0,0)$, indicating the occurrence of topological phase transitions. Since the energy gap remains open whenever the absolute value of $M_{z}$ exceeds $4\eta$ (assuming $\eta$ is positive), it is immediately evident that the system is topologically trivial in this regime. Conversely, in the range of $0<|M_{z}|<4\eta$, the gapped phases are expected to be topological. To confirm this, we numerically calculate the Chern number and determine the phase diagram shown in Fig.2(a). Within the topological region, we find that the system with a cylindrical geometry possesses one gapless chiral edge state, which aligns with the calculated Chern number, as illustrated in Fig.2(b).
Figure 2: (a) Phase diagram valid as long as $\lambda_{\rm so}\neq 0$. (b) A Chern insulator with one branch of chiral edge state. The value of the shared parameters are: $t=0$, $\lambda_{\rm so}=0.5$, $\eta=0.5$ and $M_{z}=0.3$.
Nonsymmorphic DSM in proximity to an $s$-wave superconductor.—The SOC in the DSM holds the promise of topological superconductivity if superconductivity is successfully induced into the DSM. In this paper, we explore the scenario where the DSM is grown on the surface of a conventional fully-gapped $s$-wave superconductor [see Fig.1(a)]. Through the proximity effect, the DSM inherits $s$-wave pairing from the underlying superconductor. To ensure a strong proximity effect, we reintroduce the term $\epsilon_{0}(\boldsymbol{k})$ and assume the presence of sizable Fermi surfaces in the normal state of the DSM. It is important to note that, although the upper and lower bands of the Hamiltonian (2) are separated by a gap induced by the Zeeman field, Fermi surfaces can either be generated through doping or intrinsically exist if the parameter $t$ is much larger than all other parameters.
The resulting superconducting state of the DSM can be effectively described by a Bogoliubov-de Gennes (BdG) Hamiltonian of the form [70]
@@
In the absence of a Zeeman field, the BdG Hamiltonian belongs to the DIII class according to the tenfold-way classification [72, 73]. Due to the momentum-independent nature of $s$-wave pairing, the superconducting phase is always topologically trivial in this case. However, upon breaking time-reversal symmetry through the application of a Zeeman field, the Hamiltonian transitions to the D class, thereby enabling the existence of CTSCs [70]. Regardless of the presence or absence of the Zeeman field, the BdG Hamiltonian possesses both inversion symmetry and particle-hole symmetry ($\Xi=\tau_{x}\mathcal{K}$), resulting in a symmetric energy spectrum centered at $E=0$. These two symmetries are crucial for the topological protection of the BFSs discussed later.
To demonstrate that the interplay of a Dirac band structure, a Zeeman field, and $s$-wave pairing can give rise to the emergence of CTSCs, we first consider the case where $\eta=0$. In this limit, the two sublattices are decoupled, yielding $\mathcal{H}(\boldsymbol{k})=\mathcal{H}_{A}(\boldsymbol{k})\oplus\mathcal{H}_{B}(\boldsymbol{k})$, where
@@
$\pm 2\lambda_{\rm so}(\sin k_{x}\tau_{z}s_{y}-\sin k_{y}s_{x})+\Delta_{s}\tau_{y}s_{y}.$
The SOC simplifies to the Rashba form, and this particular form of the Hamiltonians $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$ is well-known for supporting CTSCs [23, 24]. Despite the fact that the SOC terms have opposite signs for $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, their Chern numbers are identical. To facilitate a clear discussion, we assume that $t$, $M_{z}$, and $\Delta_{s}$ are all positive, with $\Delta_{s}<M_{z}$. By only varying the chemical potential $\mu$, we determine the dependence of the Chern number on $\mu$ [74], which is given by
@@
where $\mu_{c}=\sqrt{M_{z}^{2}-\Delta_{s}^{2}}$. Since $C_{A}$ equals $C_{B}$, the Chern number of the resulting CTSCs, given by $C=C_{A}+C_{B}$, is always even. Remarkably, the maximum value of $C$, and consequently the maximum number of chiral Majorana edge states, can reach as high as $4$.
@@
Intriguingly, we find that when $\eta$ attains a critical threshold, the bulk energy gap closes. However, further increasing $\eta$ does not reopen the gap. Instead, the middle two bands remain in contact, forming line degeneracies. Notably, despite the absence of a bulk gap resulting in an undefined Chern number, the chiral Majorana edge states persist if their spectral crossing occurs at $k_{y}=\pi$ [or $k_{x}=\pi$ if open (periodic) boundary conditions are considered in the $y$ ($x$) direction, due to $C_{4z}$ rotation symmetry] of the edge Brillouin zone as illustrated in Figs.3(c) and 3(d). The robustness of these chiral Majorana edge states can be attributed to the vanishing of the $\epsilon_{1}(\boldsymbol{k})$ term when $k_{x}$ or $k_{y}$ equals $\pi$. In simpler terms, when the spectral crossing of these states occurs at $k_{x}=\pi$ or $k_{y}=\pi$, increasing $\eta$ has no influence on the edge gap or the crossing at that momentum, as $\epsilon_{1}(\boldsymbol{k})$ identically vanishes at those points. In contrast, if the spectrum crossing occurs at $k_{x}=0$ or $k_{y}=0$, increasing $\eta$ causes both the edge gap and the spectrum crossing at those momenta to disappear, as illustrated in Fig.3(d).
Figure 3: Energy spectrum for a system with open (periodic) boundary conditions in the $x$ ($y$) direction. (a) TSC with $C=2$. (b) TSC with C=4C=-4. Figures (c) and (d) show the coexistence of BFS and chiral Majorana edge states. The parameters $(\mu,\eta)$ are set to $(4,0.1)$ in (a), $(0,0.025)$ in (b), $(4,0.8)$ in (c) and $(0,0.2)$ in (d), respectively. The values of the shared parameters are: $t=1$, $\lambda_{\rm so}=0.5$, $M_{z}=0.5$ and $\Delta_{s}=0.2$.
In two dimensions, line degeneracies in a superconductor correspond to topologically protected BFSs. As previously mentioned, the topological protection originates from the coexistence of inversion symmetry and particle-hole symmetry in this system. As first noted in Ref. [56], when these two symmetries coexist, the Hamiltonian can be transformed into an antisymmetric form, $\tilde{\mathcal{H}}(\boldsymbol{k})=U\mathcal{H}(\boldsymbol{k})U^{{\dagger}}$, where $U$ denotes a unitary matrix. The transformed Hamiltonian obeys the relation $\tilde{\mathcal{H}}(\boldsymbol{k})=-\tilde{\mathcal{H}}^{\mathbb{T}}(\boldsymbol{k})$, with $\mathbb{T}$ denoting the transpose operation. The antisymmetric property allows for a definition of Pfaffian, $P(\boldsymbol{k})=\text{Pf}[\tilde{\mathcal{H}}(\boldsymbol{k})]$. Since $\det\mathcal{H}(\boldsymbol{k})=\det\tilde{\mathcal{H}}(\boldsymbol{k})=P^{2}(\boldsymbol{k})$, the zeros of $P(\boldsymbol{k})$ determines the band degeneracies of $\mathcal{H}(\boldsymbol{k})$ at $E=0$. Topologically protected BFSs correspond to line degeneracies across which the sign of $P(\boldsymbol{k})$, denoted as $\text{sgn}[P(\boldsymbol{k})]$, changes from $1$ to 1-1, or vice versa.
For the current Hamiltonian, $P(\boldsymbol{k})$ is given by (see Supplemental Material (SM) [75])